# Srinivasan Kidambi, and Pravin Nair

### v1

@@ III-A2 Image super-resolution
Table III compares SwinIR and its PnP-Nystra variant on Set5 and BSDS100 across different scales for a window size of $32$ i.e. $N=32^{2}$. PnP-Nystra consistently achieves 23×2-3$\times$2 - 3 × speed-up on both GPU and CPU, with better runtime benefits at lower resolutions. The accuracy drop, measured in PSNR and SSIM, is low considering the speed-up obtained. We also demonstrate in Fig. 3 that PnP-Nystra achieves super-resolution results visually on par with the original SwinIR, with negligible perceptual difference.
III-A3 Image deblurring
Table IV reports the performance of PnP-Nystra on the RealBlur R dataset using Uformer-B for image deblurring, where window-size is $32$, i.e., $N=32^{2}$. Our method achieves a GPU speed-up of 1.76×1.76$\times$1.76 × and a CPU speed-up of 1.87×1.87$\times$1.87 × compared to the original model, with no degradation in performance (negligible change in PSNR and SSIM values). This shows that there exist cases where PnP-Nystra provides near-to-exact reconstruction as the original pre-trained models.
TABLE IV: Comparison of Uformer-B and Its PnP-Nystra Variant for Image Deblurring
Timing Analysis (GPU/ CPU, in ms )
Resolution Original PnP-Nystra Speedup
$768\times 768$ $305.16$ / $7920$ $173.41$ / $4220$ 1.76×1.76$\times$1.76 × / 1.87×1.87$\times$1.87 ×
Accuracy Analysis (PSNR / SSIM)
Dataset Original PnP-Nystra Accuracy Drop
RealBlur R $33.98$ / $0.9404$ $34.00$ / $0.9404$-0.02 dB / 0.0000
III-A4 Video super-resolution
Table V shows the runtime and accuracy of PnP-Nystra when integrated into RVRT for 4×4$\times$4 × video super-resolution, where $N$ is $64^{2}$, i.e., the window size is $64$. Compared to the original model, our method achieves over 4×4$\times$4 × inference time reduction on REDS4 and Vid4 datasets, with minimal degradation of $1.3$-$1.9$ dB in PSNR and $0.04$-$0.05$ in SSIM. Thus PnP-Nystra is an effective accelerating strategy for high-resolution video restoration. Due to hardware-specific compilation issues, we could not execute RVRT and its PnP-Nystra variant in the CPU.
TABLE V: Compariason of RVRT and its PnP-Nystra variant for ×$\times 4$ Video Super-Resolution
Timing Analysis (GPU, in ms) Accuracy Analysis (PSNR / SSIM)
Dataset Original PnP-Nystra Speedup Original PnP-Nystra Drop
REDS4 $849.53$ $207.77$ $4.09$ $30.83$ / $0.8745$ $28.95$ / $0.8338$ $1.88$ / $0.0407$
Vid4 $850.43$ $208.44$ $4.08$ $26.22$ / $0.8292$ $24.84$ / $0.7835$ $1.38$ / $0.0457$
III-B Ablation Study of PnP-Nystra
@@
In this section, we present an ablation study of PnP-Nystra, with emphasis on the internal hyper-parameters.
Approximation Error: We compute the approximation error between the attention matrix $\mathbf{S}$ of the original model and its PnP-Nystra variant, averaged over heads and layers. Note that $\mathbf{S}$ is not explicitly computed during our inference. The average mean square error between the normalized attention matrix $\mathbf{S}$ from the original model and PnP-Nystra are recorded in Table  VI. Note that elements in the matrix are in the range 00 to $1$. PnP-Nystra achieves consistently low approximation errors in the range $10^{-5}$ $10^{-6}$, across various restoration models.
TABLE VI: Attention Matrix MSE Errors for Window Attention Approximation via PnP-Nystra
Application Dataset Scale MSE Error
SwinIR image super-resolution Set5 ×$\times 2$ $1.19\times 10^{-5}$
×$\times 4$ $1.05\times 10^{-5}$
×$\times 8$ $1.24\times 10^{-5}$
BSDS100 ×$\times 2$ $1.14\times 10^{-5}$
×$\times 4$ $1.04\times 10^{-5}$
×$\times 8$ $1.32\times 10^{-5}$
Uformer image denoising SIDD $3.78\times 10^{-5}$
BSDS200 $2.00\times 10^{-5}$
Uformer image deblurring RealBlur R $6.11\times 10^{-6}$
RVRT video super-resolution REDS4 $3.89\times 10^{-5}$
Vid4 $3.00\times 10^{-3}$
Attention map comparison: Since PnP-Nystra approximates the kernel matrix $\mathbf{G}$, it implicitly approximates the attention map $\mathbf{S}$. Fig. 4 shows that PnP-Nystra yields attention patterns closely matching those of the original models.
Figure 4: Comparison of attention maps for original model and PnP-Nystra for SwinIR (top) and Uformer-B (bottom).
@@
Figure 5: Inference time vs. token count $N$: Unlike MHSA which grows quadratically with $N$, PnP-Nystra scales linearly.
Landmarks and Pseudoinverse iterations: We evaluate the effect of varying the number of landmarks $m$ and the number of iterations used for pseudoinverse computation in Step $4$ of Algorithm 1. Table VII reports PSNR and SSIM drop for the PnP-Nystra variant of Uformer-B for denoising on the SIDD dataset. As expected from Nyström approximation, the approximation performance tends to the original with an increase in number of landmarks. Additionally, a better estimate of the pseudoinverse leads to improved approximation.
TABLE VII: Impact of Number of Landmarks and Pseudoinverse Iterations on Approximation Accuracy
Varying # Landmarks (at 6 Iterations) Varying Iterations (at 32 Landmarks)
@@
$8$ $1.43$ $0.0127$ $1$ $2.38$ $0.0164$
$16$ $1.23$ $0.0119$ $3$ $1.57$ $0.0137$
$32$ $1.16$ $0.0114$ $6$ $1.16$ $0.0114$
IV Conclusion and Future Work
We presented PnP-Nystra, a plug-and-play Nyström approximation for self-attention that enables efficient inference in window-based transformer models for image and video restoration. Experiments show that PnP-Nystra acts as a fast, training-free replacement for self-attention in pre-trained models, offering a viable solution for resource-constrained deployment. Future work includes improving the speed-up of PnP-Nystra and extending the framework to global attention in vision transformers and diffusion transformers.

### v2

@@ 3.1.3 Super-resolution
Table 5 compares SwinIR and its PnP-Nystra variant on Set5 and BSDS100 across different scales for a window size of $32$, i.e., $N=32^{2}$. As shown in the Table. 2, PnP-Nystra consistently achieves $2$-$3\times$ speedup on both GPU and CPU, with better runtime benefits at lower resolutions. Even with the speedup, the accuracy drop, measured in PSNR and SSIM, is lowest among the compared linear attention methods in most evaluations in the Table. 5 across different scaling factors.
Table 5: SwinIR super-resolution accuracy (PSNR(dB),SSIM) comparison across scaling factors.
Dataset Scale Original Performer Nyströmformer PnP-Nystra
Set5 22 $36.13$ / $0.9455$ $33.91$ / $0.9322$ $35.15$ / $0.9413$ 35.25 / 0.9416
44 $30.86$ / $0.8729$ $28.11$ / $0.8142$ 29.26 / 0.8467 29.10{29.10} / 0.8437{0.8437}
88 $25.62$ / $0.7439$ $23.53$ / $0.6434$ $24.26$ / $0.6787$ 24.41 / 0.6853
BSDS100 22 $30.98$ / $0.8922$ $29.82$ / $0.8740$ $30.64$ / $0.8821$ 30.71 / 0.8830
44 $26.43$ / $0.7183$ $24.98$ / $0.6815$ $25.81$ / $0.6992$ 25.92 / 0.7008
88 $23.61$ / $0.5680$ $22.73$ / $0.5291$ $23.21$ / $0.5487$ 23.29 / 0.5503
3.2 Ablation Study of PnP-Nystra
@@
In this section, we present an ablation study of PnP-Nystra, with emphasis on the internal hyperparameters.
@@
Fig. 3: Inference time vs. token count NN: Unlike MHSA which grows quadratically with NN, PnP-Nystra scales linearly.
Landmarks and Pseudoinverse iterations: We evaluate the effect of varying the number of landmarks $m$ and the number of iterations used for pseudoinverse computation in Step $4$ of Algorithm 1. Table 6 reports PSNR and SSIM drop for the PnP-Nystra variant of Uformer-B for denoising on the SIDD dataset. As expected from Nyström approximation, the approximation performance tends to the original with an increase in the number of landmarks. Additionally, a better estimate of the pseudoinverse leads to improved approximation.
Table 6: Impact of Number of Landmarks and Pseudoinverse Iterations on Approximation Accuracy
Varying # Landmarks (at 6 Iterations) Varying Iterations (at 16 Landmarks)
@@
88 $1.54$ $0.0127$ 11 $2.84$ $0.0167$
$16$ $1.01$ $0.0108$ 33 $1.94$ $0.0138$
$32$ $0.86$ $0.0102$ 66 $1.01$ $0.0108$
Attention map comparison: PnP-Nystra approximates the kernel matrix $\mathbf{G}$ and therefore induces an approximate attention map $\mathbf{S}$. As shown in Fig. 4, the resulting attention distribution closely follows that of the original model, highlighting the same prominent edges and structural regions while maintaining similar suppression over less informative areas.
Fig. 4: Attention map visualization for Uformer-B: (a) original model and (b) PnP-Nystra. Both maps have the same salient structures with strong responses along object boundaries
4 Conclusion and Future Work
We presented PnP-Nystra, a plug-and-play Nyström approximation for self-attention that enables efficient inference in window-based transformer models for image restoration. Experiments show that PnP-Nystra acts as a fast, training-free replacement for self-attention in pre-trained models, offering a viable solution for resource-constrained deployment. Future work will focus on extending PnP-Nystra to global attention in vision transformers and diffusion transformers, and spatiotemporal attention for video restoration.